# The mean solar butterfly diagram and poloidal field generation rate at the surface of the Sun

### v1

@@ Abstract
Methods. We use data from Wilcox Solar Observatory to generate time-latitude diagrams spanning cycles 21 to 24 of the surface radial and toroidal magnetic fields, symmetrize them across the equator and cycle-average them. From the mean butterfly diagram and surface toroidal field we then infer the mean poloidal field generation rate at the surface of the Sun.
@@
Conclusions. Cycle-averaging butterfly diagrams removes realization noise and artefacts due to imperfect scale separation, and corresponds to an ensemble average that can be interpreted in the mean-field framework. The result can then be directly compared to $\alpha\Omega$-type dynamo models. The Babcock-Leighton $\alpha$-effect is consistent with observations, a result that can be appreciated only if the observational data is averaged in some way.
@@
Sun: magnetic fields – Sun: activity
@@ 1 Introduction
The solar cycle lasts on average approximately 11 years and is believed to be driven by a hydromagnetic dynamo seated in the convection zone of the Sun (Charbonneau 2020). The dynamo loop consists of two parts; one where toroidal field is generated from the poloidal field, and another where poloidal field of opposite sign is generated from the new toroidal field. It takes another solar cycle, or half of the magnetic cycle, to revert to the original polarity. The first part is relatively well understood and takes place through the so-called $\Omega$-effect, whereby poloidal field is wound up in the azimuthal direction by the differential rotation, which is observationally well constrained by helioseismology (eg. Schou et al. 1998). The second part, on the other hand, involves the unobservable subsurface toroidal field and two possible mechanisms; the turbulent $\alpha$-effect (Parker 1955a; Steenbeck et al. 1966) and the Babcock-Leighton (BL) mechanism (Babcock 1961; Leighton 1964, 1969). Since the latter can also be put in a functional form identital to the former, it is sometimes called the BL $\alpha$-effect.
@@
The mean field of mean-field electrodynamics is defined in terms of an average that can either be spatial, temporal or ensemble. In the case of the Sun, the azimuthal average is a natural choice. In fact, the butterfly diagram is constructed by longitudinally averaging line-of-sight synoptic magnetograms. The symmetric component (with respect with the solar meridian) is then extracted, from which the radial field is obtained by assuming it is the only component of the poloidal field at the surface. Finally, the butterfly diagram is obtained by stacking the result in time. In principle, butterfly diagrams are thus of mean-field nature and can be directly compared to models.
@@
A caveat, however, is that scale separation is likely to be poor in the Sun as supergranulation reaches down to about 5% of the solar radius, which is non-negligible with respect to the size of the convection zone and the scales of the differential rotation and meridional flow. The picture is much worse if we consider the scale of active regions, which is on the order of 100 Mm, about half the depth of the convection zone. Consequently the observed butterfly diagram can’t be directly compared to simple mean-field models aimed at reproducing the most fundamental features of the solar cycle. A way to circumvent this is by considering different solar cycles as different realizations of the dynamo, and ensemble average them (note that this is not a true ensemble average in the mean-field sense, cf. Hoyng 2003, but is aimed at reproducing the azimuthal average used in models with reduced realization noise). In other words, average different cycles together to remove realization noise (fluctuating dynamo parameters) and the effects of imperfect scale separation.
@@
The way cycles are to be averaged together requires careful consideration. It has been shown that defining cycles in terms of activity minimum or maximum is not ideal (Waldmeier 1955; Hathaway 2011). Nevertheless, observations show that all cycles present the same equatorward pattern of sunspot migration, regardless of strength (Waldmeier 1939; Hathaway 2011). When cycles are defined according to the time the centroid of sunspot area reach a certain latitude, they all behave the same way; the location of this centroid migrates equatorward according to the same standard law (Hathaway 2011). For this reason, we define the reference times as 4 years before the times when the centroid of sunspot area reach $19^{\circ}$ (Cameron & Schüssler 2023). The individual cycles are then averaged together in phase. To carry out this cycle-averaging we use butterfly diagrams produced from synoptic magnetograms obtained by the Wilcox Solar Observatory (WSO). We further symmetrize the data across the equator so that we have, in effect, an average consisting of 8 cycles.
@@
In addition to the surface radial field, another important quantity we can infer from longitudinally-averaged line-of-sight magnetograms is the toroidal field emerging at the surface, the antisymmetric component (with respect to the solar meridian), from which we can construct a time-latitude diagram (from WSO data, Cameron et al. 2018). Having both mean time-latitude diagrams of the surface radial and toroidal fields at our disposal, we can test the validity of the simple $\alpha\Omega$ dynamo models, as well as possibly constrain the functional form of the source term. To do this, we calculate independently from both quantities an intermediate one: the surface radial field generation rate – the radial source term. This quantity can be inferred from the radial field butterfly diagram using the surface flux transport (SFT) model (Leighton 1964; DeVore et al. 1984; Sheeley et al. 1985), and independently from the surface toroidal field butterfly diagram using Joy’s law. Both diagrams are then compared to determine if they are consistent with each other.
Figure 1: Time-latitude diagrams of the surface radial field $B_{r}$ (top) and of the surface toroidal field $B_{\phi}$ (bottom). The vertical dotted lines indicate the cycle start times as given by the times where the central latitude of the sunspot belts reach $19^{\circ}$ minus 4 years.
Figure 2: Time-latitude diagrams of the cycle-averaged surface radial field $B_{r}$ (top) and of the surface toroidal field $B_{\phi}$ (bottom). The vertical dotted lines indicate the cycle start times as given by the times where the central latitude of the sunspot belts reach $19^{\circ}$ minus 4 years.
@@ 2 The observed mean butterfly diagram
The WSO data we use is presented in Figure 1. Also shown are the reference start times of the individual cycles, defined by the time when the central latitude of the activity belts reach $19^{\circ}$ minus 4 years (Cameron & Schüssler 2023).
We have 4 cycles of data for 2 hemispheres, so that we have in effect 8 cycles to average together. We perform the cycle-averaging as follows: we first symmetrize the data across the equator, then for each cycle shift the butterfly diagram starting from that cycle backwards to the reference time of cycle 21, and average. The result of this procedure is shown in Figure 2. It consists of 4 averages of the solar cycles obtained by a decreasing amount of cycles.
One can immediately see that the cycle-averaging produces butterfly wings that are much better defined. We have clear butterfly wings of one polarity, that of the leading sunspots, and poleward migrating fields of the opposite polarity, that of the trailing sunspots. The latter feature is also known as the ”rush to the poles” (Ananthakrishnan 1954). Even with only 8 realizations, the trailing polarity field is essentially removed from the butterfly wings. The rush to the poles is also increasingly filled with fields of trailing polarity and devoid of fields of leading polarity. We thus see a ”separation of polarities” happening. The butterfly wings are much more clearly delineated and seem narrower with a width slightly under $30^{\circ}$, rather than one of $\simeq 35^{\circ}$. Lastly, the rapid increase of sunspot number during the rising phase, and its slow decrease during the declining phase is much more apparent.
It is clear, however, that 4 solar cycles is not nearly enough to carry out a proper ensemble average. It is nonetheless reasonable, from these results, to assume that the addition of more cycles would further the observed trend. Most of the small-scale features will be washed out and the mean observed butterfly diagram will look very similar to those calculated from mean-field dynamo models; smooth butterfly wings of the leading spot polarity, and rush to the poles and polar fields of trailing spot polarity.
Note the change of the scaling from Figure 1 to 2. The largest fields inside the butterfly wings are brought down from 3 G to 2 G. Higher-resolution butterfly diagrams (eg. Norton et al. 2023) have fields the order of 10 G; in WSO data part of the averaging has already been done by the instrument’s limited resolution. This is more evidence that one needs to be careful when comparing simulation results from mean-field models to observations.
@@ 3 Inferring the radial field generation rate
The radial field generation rate, or the radial source term, can be inferred both through the butterfly diagram and the time-latitude diagram of the surface toroidal field, by making use of models of surface flux transport (SFT) and emergence (BL mechanism).
@@ 3.1 Surface flux transport model
The SFT model describes the passive advective and diffusive transport of the surface radial field by the following equation (Yeates et al. 2023):
@@
where $B_{r}$, $\omega(\lambda)$, $v(\lambda)$, $\eta$, and $S_{r}(\lambda,\phi,t)$ are respectively the surface radial field, angular velocity, meridional flow, diffusivity, and radial source term, $\lambda$ is the latitude and $\phi$ the longitude. Since the magnetic butterfly diagram provides us the longitudinally-averaged surface radial field $B_{r}(\lambda,t)$, the only unknown is the surface radial source term
@@
We use a value of turbulent diffusivity of $\eta=350\leavevmode\nobreak$km2/s, which is consistent with estimates from observations (eg. Komm et al. 1995), the SFT model of Lemerle et al. (2015) and MLT (eg. Muñoz-Jaramillo et al. 2011). As for the surface meridional flow profile, we use the anti-symmetrized time-dependent profiles obtained from the helioseismic inversions of Gizon et al. (2020).
@@ 3.2 Emergence model
There is strong evidence (see eg. Dasi-Espuig et al. 2010; Kitchatinov & Olemskoy 2011; Cameron & Schüssler 2015) that suggests the BL mechanism is the main poloidal field generation mechanism of the solar dynamo. The first part of this mechanism is the emergence through the photosphere of tilted toroidal flux tubes. The corresponding generation of (radial) poloidal field from the emerging toroidal field can be derived as follows.
We begin with the solenoidal condition $\nabla\cdot{\boldsymbol{B}}=0$ for an axisymmetric field ${\boldsymbol{B}}$,
@@
Joy’s law implies that as flux is emerging through the photosphere $B_{\theta}=B_{\phi}\tan\delta$, where $\delta$ is the tilt angle of the resulting bipolar magnetic region. The rate at which the flux of $B_{\phi}$ is carried through the surface per radian in $\theta$ is
@@
where $v_{e}$ is the radial velocity of the emerging field. Then
@@
We use the form of Joy’s law as given by Leighton (1969), $\sin\delta=0.5\sin\lambda$. As for the emergence velocity, we consider three different prescriptions. The first is the constant velocity of $v_{e}\simeq 200\leavevmode\nobreak$m/s inferred by Centeno (2012). We also consider the cases where the emergence velocity $v_{e}$ is proportional to the toroidal field $B_{\phi}$ (Parker 1955b) or to its square $B_{\phi}^{2}$ (Unno & Ribes 1976). We thus write the emergence velocity as
@@
Figure 3: Left panel: time-latitude diagrams of the surface radial source term $S_{r}$ inferred from the high-res (top) and low-res (middle) data by use of the SFT model, and from the low-res surface toroidal field (bottom). The vertical dotted lines indicate the cycle start times as given by the times where the central latitude of the sunspot belts reach $19^{\circ}$ minus 4 years.
Figure 4: Time-latitude diagrams of the surface source term as obtained from the SFT model (upper left) and emergence model with different emergence velocity prescriptions; constant velocity (lower left), linear (upper right) and quadratic (lower right) toroidal field dependence on the emergence velocity.
@@ 4 The observed mean radial field generation rate
The mean surface radial field generation rates $S_{r}$, as inferred from the surface radial and toroidal fields, are shown in Figure 3. We clearly see one ”mean BMR” inside the activity belts. Remarkably, both methods used to infer the radial field generation rate give results in good qualitative agreement, including both the shape of the mean BMRs and their amplitudes. There are, however, three main noticeable discrepancies. Firstly, the SFT model gives us important radial field generation at the poles, of the opposite polarity to that of the polar fields. This could be explained by underestimated polar field strengths from the observations (Linker et al. 2017), possibly in combination with a weakened meridional flow amplitude at latitudes above $60^{\circ}$ (Mahajan et al. 2021). Secondly, there are sorts of weak amplitude ”bridges” at cycle minimum between the poleward polarity of one cycle and the equatorward polarity of the next (showcasing the magnetic cycle and activity overlap between cycles). This overlap is only visible in the $S_{r}$ obtained through the surface toroidal field and Joy’s law. Thirdly, the source term derived from the surface toroidal field and Joy’s law are slightly wider and are shifted slightly polewards compared to those obtained from the surface radial field.
@@
Lastly, we consider emergence velocities of the toroidal field depending linearly and quadratically on the toroidal field strength. The result is presented in Figure 4. We see that increasing this dependence makes the mean BMR restricted to increasingly lower latitudes. Furthermore, in both the linear and quadratic cases, the aforementioned ”bridges” between sunspot cycles seen in the constant velocity case disappear. Since the bridges are lacking in the source term necessary to explain the evolution of the surface radial field, this suggests that either Joy’s law or the emergence velocity of low flux regions might differ from that of active regions.
5 Conclusion

### v2

@@ Abstract
Methods. We used data from Wilcox Solar Observatory to generate time-latitude diagrams of the surface radial and toroidal magnetic fields spanning cycles 21 to 24. We symmetrized them across the equator and cycle-averaged them. From the mean butterfly diagram and surface toroidal field, we then inferred the mean poloidal field generation rate at the surface of the Sun.
@@
Conclusions. Cycle-averaging butterfly diagrams removes realization noise and artefacts due to imperfect scale separation and corresponds to an ensemble average that can be interpreted in the mean-field framework. The result can then be directly compared to $\alpha\Omega$-type dynamo models. The Babcock-Leighton $\alpha$-effect is consistent with observations, a result that can be appreciated only if the observational data are averaged in some way.
@@
Sun: magnetic fields – Sun: activity – Sun: interior
@@ 1 Introduction
The 11-year solar cycle is believed to be driven by a hydromagnetic dynamo seated somewhere in the convection zone of the Sun (Charbonneau 2020). The dynamo loop consists of two parts: one where toroidal field is generated from the poloidal field, and the other where poloidal field of opposite sign is generated from the new toroidal field. It takes another solar cycle, or half of the magnetic cycle, to revert to the original polarity. The first part is relatively well understood and takes place via the $\Omega$-effect, whereby the poloidal field is wound up in the azimuthal direction by differential rotation, which is observationally well constrained by helioseismology (e.g. Schou et al. 1998). The second part, on the other hand, involves the unobservable subsurface toroidal field and two possible mechanisms: the $\alpha$-effect (Parker 1955b; Steenbeck et al. 1966) and the Babcock-Leighton (BL) mechanism (Babcock 1961; Leighton 1964, 1969). This dynamo loop is often studied using mean-field models.
@@
The mean field of mean-field electrodynamics is defined in terms of an average that can be spatial, temporal, or ensemble. In the case of the Sun, the azimuthal average is a natural choice. In fact, the butterfly diagram is constructed by longitudinally averaging line-of-sight synoptic magnetograms. The symmetric component (with respect to the solar meridian) is then extracted, from which the radial field is obtained by assuming it is the only component of the poloidal field at the surface. Finally, the butterfly diagram is obtained by stacking the result in time. In principle, butterfly diagrams are thus of a mean-field nature and can be directly compared to models.
@@
A caveat, however, is that scale separation is likely to be poor in the Sun as supergranulation reaches down to about 5% of the solar radius, which is non-negligible with respect to the depth of the convection zone and the scales of differential rotation and the meridional flow. The picture is much worse if we consider active regions (which, by being largely non-axisymmetric, are supposed to be small-scale in nature), with sizes of the order of 100 Mm, or about half the depth of the convection zone. Consequently, the observed butterfly diagram cannot be directly compared to mean-field models. A way to circumvent this problem is by considering different solar cycles as different realizations of the underlying dynamo and ensemble-averaging them; we note that this is not a true ensemble average in the mean-field sense (see Hoyng 2003) but is aimed at reproducing the azimuthal average used in models with reduced realization noise. In other words, averaging different cycles together to remove realization noise (fluctuating dynamo parameters) and the effects of imperfect scale separation.
Figure 1: 12-month average of the sunspot area (top) and centroid latitudes as a function of time starting from $t_{0}$ (bottom). The dashed green lines represent the centroid latitude of the sunspot zone (Eq. (1)) and the vertical dashed lines the reference start times of the cycles (see text).
@@
The way cycles are to be averaged together requires careful consideration. It has been shown that defining cycles in terms of activity minimum or maximum is not ideal (Waldmeier 1935; Hathaway 2011). But when cycles are instead defined according to their start time, they all present the same equatorward pattern of sunspot migration, regardless of strength; the location of the sunspot area centroid migrates towards the equator according to the same standard law (Waldmeier 1939, 1955; Hathaway 2011; Cameron & Schüssler 2023). The individual cycles can then be averaged together in phase. To carry out this cycle-averaging, we used butterfly diagrams produced from synoptic magnetograms obtained at Wilcox Solar Observatory (WSO). We further symmetrized the data across the equator so that we have, in effect, an average consisting of eight cycles.
@@
In addition to the surface radial field, another important quantity we can infer from longitudinally averaged line-of-sight magnetograms is the toroidal field emerging at the surface, the antisymmetric component (with respect to the solar meridian), from which we can construct a time-latitude diagram (from WSO data; Cameron et al. 2018). Having mean time-latitude diagrams of both the surface radial and toroidal fields at our disposal, we can test the validity of the simple $\alpha\Omega$ dynamo models, as well as possibly constrain the functional form of the source term. To do this, we used both quantities to independently calculate an intermediate one: the surface radial field generation rate (i.e. the radial source term). We inferred this quantity from the butterfly diagram using the surface flux transport (SFT) model (Leighton 1964; DeVore et al. 1984; Sheeley et al. 1985), and from the time-latitude diagram of the surface toroidal field using Joy’s law. We then compared the two diagrams to determine if they are consistent with each other.
Figure 2: Cycle average of the sunspot area (top) and centroid latitudes as a function of time starting from $t_{0}$ (bottom). The dashed green lines represent the centroid latitude of the sunspot zone (Eq. (1)) and the vertical dashed lines the reference start times of the cycles (see the main text).
2 The equatorward migration of the sunspot zones
The original work of Waldmeier (1939, 1955) concerning the equatorward migration of the sunspot zones was extended by Hathaway (2011), who used the then entire daily sunspot area record (cycles 12 to 22) as provided by the Royal Observatory, Greenwich (May 1874 to December 1976) and the United States Air Force and National Oceanic and Atmospheric Administration (from January 1977). The average daily sunspot area for each Carrington rotation, $A(\lambda_{i})$, was then calculated, as was the latitude of the centroid of the sunspot area for each hemisphere and Carrington rotation, separately for each cycle, according to a sunspot area weighted average,
$\lambda_{\mathrm{c}}=\frac{\sum_{i}A(\lambda_{i})\lambda_{i}}{\sum_{i}A(\lambda_{i})},$ (1)
the sums being carried out over the 25 latitude bins of each hemisphere. The start time of each cycle, $t_{0}$, was then determined by fitting the monthly sunspot number record of each cycle to a skewed Gaussian devised by Hathaway et al. (1994). This procedure allowed Hathaway (2011) to confirm the results of Waldmeier (1939, 1955), that the latitudinal drift of the sunspot zones follows the same universal path, regardless of cycle strength or phase. This standard law was found to closely approximate the following decaying exponential:
$\lambda_{\mathrm{c}}(t)=28^{\circ}\exp{\left(-\frac{t-t_{0}}{90\leavevmode\nobreak\ \textrm{mths}}\right)}.$ (2)
Our cycle-averaging procedure requires the determination of $t_{0}$ for each cycle. To do so, we first took the 12-month average of the sunspot area, only considering the four cycles for which we have magnetograms (Fig. 1). [^1] Downloaded from http://solarcyclescience.com/activeregions.html. We then determined the centroid latitudes of the sunspot belts by applying Eq. (1), averaged both hemispheres together, and lastly fitted the first six years of these curves with Eq. (2). The result is shown in Fig. 1.
We considered the average of the quantity $f$ defined as
$\overline{f}(\theta,t-t_{01})=\sum_{i=1}^{n}\frac{f(\theta,t-t_{0i})}{n_{k}(t)},$ (3)
$n$ being the number of cycles that are averaged and $n_{k}(t)=n+i-k$ for $t_{0k}\leq t<t_{0k+1}$. The cycle-averaging was performed as follows: four copies of the data were made; then for each copy, $i$, the whole dataset was shifted back so that the cycle start times $t_{0i}$ and $t_{01}$ coincide, whereafter the shifted copies were averaged together. We further symmetrized the data across the equator. Our procedure explicitly minimizes any north-south asymmetry, which will also contain information about the dynamo. The result for the 12-month average of the sunspot area is shown in Fig. 2. It consists of four averages of the solar cycles obtained by including a decreasing amount of cycles in the average.
@@ 3 The observed mean butterfly diagram
The WSO data we used are presented in Fig. 3 along with the reference start times of the individual cycles (see Sect. 2). By carrying out our cycle-averaging procedure, we obtained the time-latitude diagrams presented in Fig. 4. One can immediately see that the cycle-averaging produces butterfly wings that are much better defined. We have clear butterfly wings of one polarity, that of the leading sunspots, and poleward migrating fields of the opposite polarity, that of the trailing sunspots. The latter feature is also known as the ‘rush to the poles’ (Ananthakrishnan 1954; Altrock 1997). Even with only eight realizations, the trailing polarity field is essentially removed from the butterfly wings. The rush is also increasingly filled with fields of trailing polarity and is devoid of fields of leading polarity (see Fig. 3 of Mordvinov et al. 2022, for a more detailed look at the small-scale surges and features). We thus see a ‘separation of polarities’ happening. The butterfly wings are much more clearly delineated and seem narrower, with a width slightly under $30^{\circ}$ rather than $\simeq 35^{\circ}$. Lastly, the rapid increase in the sunspot number during the rising phase, and its slow decrease during the declining phase, is much more apparent.
Figure 3: Time-latitude diagrams of the surface radial field, $B_{r}$ (top), and surface toroidal field, $B_{\phi}$ (bottom). The vertical dotted lines indicate the cycle start times as determined by the procedure outlined in Sect. 2.
It is clear, however, that four solar cycles is not nearly enough to carry out a proper ensemble average. It is nonetheless reasonable to assume that the addition of more cycles would only further the observed trend. Most of the small-scale features will be washed out and the mean observed butterfly diagram will look very similar to those calculated from mean-field dynamo models: smooth butterfly wings of the leading spot polarity, and a rush to the poles and polar fields of the trailing spot polarity.
We note the change of the scaling from Fig. 3 to 4: the largest fields inside the butterfly wings are reduced from 3 G to 2 G. Higher-resolution butterfly diagrams (e.g. Norton et al. 2023) have fields of the order of 10 G; in WSO data, part of the averaging has already been done by the instrument’s limited resolution. This is more evidence that one needs to be careful when comparing simulation results from mean-field models to observations.
@@ 4 Inferring the radial field generation rate
The radial field generation rate, or the radial source term, can be inferred both through the butterfly diagram and the time-latitude diagram of the surface toroidal field by using SFT and emergence (BL mechanism) models.
Figure 4: Time-latitude diagrams of the mean surface radial field, $\overline{B_{r}}$ (top), and surface toroidal field, $\overline{B_{\phi}}$ (bottom). The vertical dotted lines indicate the cycle start times as determined by the procedure outlined in Sect. 2.
@@ 4.1 Surface flux transport model
The SFT model describes the passive advective and diffusive transport of the surface radial field via the following equation (Yeates et al. 2023):
@@
where $B_{r}$, $\omega(\lambda)$, $v(\lambda)$, $\eta$, and $S_{r}(\lambda,\phi,t)$ are respectively the surface radial field, angular velocity, meridional flow, diffusivity, and radial source term; $\lambda$ is the latitude and $\phi$ the longitude. Since the magnetic butterfly diagram provides us with the longitudinally averaged surface radial field, $B_{r}(\lambda,t)$, the only unknown is the surface radial source term,
@@
We used a value of turbulent diffusivity of $\eta=350\leavevmode\nobreak$km2/s, which is consistent with estimates from observations (e.g. Komm et al. 1995), the SFT model from Lemerle et al. (2015), and mixing length theory (e.g. Muñoz-Jaramillo et al. 2011). As for the surface meridional flow profile, we used the anti-symmetrized time-dependent profiles obtained from the helioseismic inversions of Gizon et al. (2020).
@@ 4.2 Emergence model
There is strong evidence (see e.g. Dasi-Espuig et al. 2010; Kitchatinov & Olemskoy 2011; Cameron & Schüssler 2015) that the BL mechanism is the main poloidal field generation mechanism of the solar dynamo. The first part of this mechanism is the emergence through the photosphere of tilted toroidal flux tubes. The corresponding generation of radial field from the emerging toroidal field is derived as follows.
Figure 5: Time-latitude diagrams of the mean surface radial source term, $\overline{S_{r}}$, inferred from the surface radial field and SFT model (top) and from the surface toroidal field and emergence model for $m=0$ (bottom); see the main text. The vertical dotted lines indicate the cycle start times as determined by the procedure outlined in Sect. 2.
Figure 6: Time-latitude diagrams of the mean surface radial source term, $\overline{S_{r}}$, inferred from the surface toroidal field and emergence model for $m=1$ (top) and $m=2$ (bottom); see the main text. The vertical dotted lines indicate the cycle start times as determined by the procedure outlined in Sect. 2.
We begin with the solenoidal condition $\nabla\cdot{\boldsymbol{B}}=0$ for an axisymmetric field, ${\boldsymbol{B}}$,
@@
Joy’s law implies that as flux is emerging through the photosphere, $B_{\theta}=B_{\phi}\tan\delta$, where $\delta$ is the tilt angle of the resulting bipolar magnetic region (BMR). The rate at which the flux of $B_{\phi}$ is carried through the surface per radian in $\theta$ is
@@
where $v_{e}$ is the radial velocity of the emerging field. Thus,
@@
We used the form of Joy’s law as given by Leighton (1969), $\sin\delta=0.5\sin\lambda$. As for the emergence velocity, we considered three different prescriptions. The first is the constant velocity of $v_{e}\simeq 200\leavevmode\nobreak$m/s inferred by Centeno (2012); we also considered the cases where the emergence velocity, $v_{e}$, is proportional to the toroidal field, $B_{\phi}$ (Parker 1975) or to its square, $B_{\phi}^{2}$ (Unno & Ribes 1976), mimicking magnetic buoyancy (Parker 1955a; Jensen 1955). We thus write the emergence velocity as
@@ 5 The observed mean radial field generation rate
The mean surface radial field generation rates, $\overline{S_{r}}$, as inferred from the surface radial (Sect. 4.1) and toroidal fields (Sect. 4.2), are shown in Fig. 5. We clearly see one ‘mean BMR’ inside the activity belts. Remarkably, the two methods used to infer the radial field generation rate give results that are in good qualitative agreement, including for both the shape of the mean BMRs and their amplitudes. There are, however, three main noticeable discrepancies. Firstly, the SFT model gives us important radial field generation at the poles, of the opposite polarity to that of the polar fields. This could be explained by underestimated polar field strengths from the observations (Linker et al. 2017), possibly in combination with a weakened meridional flow amplitude at latitudes above $60^{\circ}$ (Mahajan et al. 2021). Secondly, there are sorts of weak-amplitude ‘bridges’ at cycle minima between the poleward polarity of one cycle and the equatorward polarity of the next (showcasing the magnetic cycle and activity overlap between cycles). This overlap is only visible in the $\overline{S_{r}}$ obtained through the surface toroidal field and Joy’s law. Thirdly, the source terms derived from the surface toroidal field and Joy’s law are slightly wider and are shifted slightly polewards compared to those obtained from the surface radial field.
@@
Lastly, we assumed that the emergence velocities of the toroidal field depend linearly and quadratically on the toroidal field strength. The result is presented in Fig. 6. We note that the cycle average was carried out after determining the surface radial source term. We see that increasing this dependence restricts the mean BMR to increasingly lower latitudes. Furthermore, in both the linear and quadratic cases, the aforementioned ‘bridges’ between sunspot cycles seen in the constant velocity case disappear. Since the bridges are absent in the $\overline{S_{r}}$ time-latitude diagram necessary to explain the evolution of the surface radial field, this suggests that either Joy’s law or the emergence velocity of low flux regions might differ from that of active regions.
6 Conclusions